Find the equation of the line which passes through the point (1,√3) and makes an angle of 30◦ with the line x-√3y+√3 = 0.
1
Expert's answer
2021-11-12T08:18:09-0500

Write the equation of the line in slope-intercept form



"y=mx+b"

Line


"x-\sqrt{3}y+\sqrt{3}=0""y=\dfrac{1}{\sqrt{3}}x+1""slope_1=m_1=\dfrac{1}{\sqrt{3}}"

The angle between two lines



"\tan \theta=\pm\dfrac{m_2-m_1}{1+m_1m_2}"

Given "\theta=30\degree"



"\tan 30\degree=\pm\dfrac{m_2-m_1}{1+m_1m_2}""\dfrac{1}{\sqrt{3}}=\pm\dfrac{m_2-\dfrac{1}{\sqrt{3}}}{1+\dfrac{1}{\sqrt{3}}m_2}""\sqrt{3}+m_2=\pm(3m_2-\sqrt{3})"

Let



"\sqrt{3}+m_2=-(3m_2-\sqrt{3})""m_2=0"

The equation of the line which passes through the point "(1, \sqrt{3})" is



"y=\sqrt{3}"


Let



"\sqrt{3}+m_2=(3m_2-\sqrt{3})""m_2=\sqrt{3}"

Point "(1, \sqrt{3})"

"\sqrt{3}=\sqrt{3}(1)+b=>b=0"

The equation of the line which passes through the point "(1, \sqrt{3})" is



"y=\sqrt{3}x"

"y=\sqrt{3}" or "y=\sqrt{3}x"